OpenStudy (geekfromthefutur):

PLZ help ive tried google and it wont help I will fan you and give you a medal The table below shows some values of f(x) and g(x) for different values of x: Complete the chart and determine the solution of the equation f(x) = g(x) x = -2 x = -1 x = 1 x = 4

OpenStudy (geekfromthefutur):

x f(x) = 3x + 7 g(x) = 4)-x -2 16 -1 4 0 1 1 2 13

ganeshie8 (ganeshie8):

if possible, please take a screenshot of complete question and attach...

OpenStudy (geekfromthefutur):

It wont let me on desktop

OpenStudy (geekfromthefutur):

@ganeshie8

OpenStudy (geekfromthefutur):

@ganeshie8 i got a screen shot

ganeshie8 (ganeshie8):

when x = -1, whats the value of g(x) ?

ganeshie8 (ganeshie8):

\[\large g(x) = 4^{-x}\]

ganeshie8 (ganeshie8):

plugin x = -1, what do u get ?

ganeshie8 (ganeshie8):

you get : \[\large g(-1) = 4^{-(-1)} = 4^1 = 4\] which is same as f(-1), right ?

OpenStudy (geekfromthefutur):

right

ganeshie8 (ganeshie8):

that means, x=-1 is a solution to the given equations